Question Ref No: 3078 Vevoophonexdekedut Sdn Bhd began the year with total assets of RM586000 and total liabilities of RM266000 Required Using this information and the accounting equation, calculate for each of the following independent questions 1.What was Vevoophonexdekeduts shareholders equity at the beginning of the year? Format: 420000 2.Assuming Vevoophonexdekeduts assets increased by RM50000 and its total liabilities increased by RM33000 during the year, what would be the amount of shareholders equity at the end of the year? 3.Assuming Vevoophonexdekeduts total assets increased to RM704000 and its shareholders equity increased to RM295000, what Format: 748000 would be the amount of total liabilities at the end of the year? Format 902000
Answer #1

1.

shareholders'equity at the beginning of the year

=total assets-total liabilities

=586000-266000

=320000

2.

shareholders'equity at the end of the year

=total assets-total liabilities

=(586000+50000)-(266000+33000)

=337000

3.

total liabilities at the end of the year

=total assets-shareholders'equity

=704000-295000

=409000

the above is answer..